# Weiqi Wang

### v1

@@ C.1.4 Metaphysical Transition Generation
C.2 Main Evaluations on [Uncaptioned image]Mars
To evaluate LLMs on three tasks in [Uncaptioned image]Mars, we show our evaluating prompts in zero-shot scenario in Table 5. These models are consistent across all models’ evaluations for fair comparison. For few-shot evaluations, few shot examples are added after task descriptions and before the prompted test entry. The exemplars are randomly sampled for each different test entry. For COT prompting, we specifically ask LLMs to “think step by step and generate a short rationale to support your reasoning.” Then, we ask it to give an answer based on its generated rationale. The sampling temperature $\tau$ is set to 0.1 by default, and 5 COT responses are sampled with $\tau$ set to 0.7 in the SC-COT setting.
@@ C.2 Main Evaluations on [Uncaptioned image]Mars
Appendix D Implementation Details
@@
For all experiments, we use the Huggingface [^1] https://huggingface.co/ Library [70] to build all models. For each LLM, we conduct experiments with both its instruction fine-tuned version (if any) and the original version. The one achieving higher performances will be included in the reported results. For LLaMa2, the model code is meta-llama/Llama-2-7b/13b/70b(-chat)-hf. For LLaMa3, the model code is meta-llama/Meta-Llama-3-8B/70B(-Instruct). For Mistral, we use mistralai/Mistral-7B(-Instruct)-v0.2.
@@
All experiments are conducted on eight NVIDIA-V100 (32G) GPUs, with 8E disk space, 48 CPU cores, and 1T memory. Each experiment is repeated three times with different random seeds, and the average performances are reported.
D.1 Main Evaluations on [Uncaptioned image]Mars
@@
To evaluate PTLMs in a zero-shot manner, we adopt the evaluation pipeline used for zero-shot question answering [44, 63]. Specifically, we convert each discrimination data entry into two declarative statements, which serve as natural language assertions corresponding to ‘yes” or “no” options. For instance, when determining whether an event is metaphysical, we generate two assertions: “The event <EVENT> is metaphysical as it’s unlikely to occur in reality,” and “The event <EVENT> is not metaphysical; it’s plausible in reality.” The models are then tasked with computing the loss of each assertion. The assertion with the lowest loss is considered as the model’s prediction. This approach allows any PTLM to be evaluated under classification tasks with an arbitrary number of options or even type classification based on a single assertion. We use the open code library [^3] https://github.com/Mayer123/HyKAS-CSKG as our code base and follow the default hyperparameter settings. For VERA, we follow the exact same implementation [^4] https://github.com/liujch1998/vera [40]. The accessed backbone model is liujch1998/vera, and all other hyperparameter settings follow the default implementation.
For fine-tuning PTLMs, we connect each PTLM backbone with five fully connected classification layers. The entire model is then fine-tuned using a classification objective with cross-entropy loss. We employ a default setting of a learning rate of 5e-6 and a batch size of 64. The models are optimized using an AdamW optimizer [43], with the model’s performance evaluated every 50 steps. We set the maximum sequence lengths for the tokenizers to 70 for all three discriminative subtasks. Early stopping is also implemented to select the best checkpoint when the highest validation accuracy is achieved. To ensure convergence, we train all models with five epochs.
@@
For fine-tuning LLMs, we use LoRA for fine-tuning, and the LoRA rank and $\alpha$ are set to 16 and 32, respectively. We adopt the open code library from LlamaFactory [^5] https://github.com/hiyouga/LLaMA-Factory [78] for model training and evaluation. We similarly use an Adam [36] optimizer with a learning rate of 5e-5 and a batch size of 8. The maximum sequence length for the tokenizer is set at 300. All models are fine-tuned over three epochs, selecting the checkpoint with the highest accuracy on the validation set.
Finally, for evaluating proprietary LLMs, such as ChatGPT and GPT4, we similarly prompt them as with open LLMs. Detailed prompts are explained in Appendix C.2.
D.2 Improving Metaphysical Reasoning via Transferring from Conceptualization Taxonomy
@@
CANDLE’s data is primarily divided into two sections. The first section comprises conceptualizations of instances or events, which can be reformatted into metaphysical event discrimination. Each data entry in CANDLE represents a conceptualization of an abstracted instance within an event or the abstraction of an entire event. Following our definition in Section 3, we interpret each conceptualization as a change in the event. For each data entry, replacing the original instance with its conceptualization forms a plausible change that could occur in reality. Subsequently, we randomly select negative conceptualizations for an event from conceptualizations of other events that do not share any common words with the anchor event. These negative conceptualizations form metaphysical events. Three models are then pre-trained on four million events, with a balanced ratio of plausible events and metaphysical events. The hyperparameters for fine-tuning all models remain consistent with the implementation details described above in Appendix D.1.
@@
Appendix E Annotation Details
E.1 Worker Selection Protocol
@@
E.2 Annotation Interface
For each task, we provide workers with comprehensive task explanations in layman’s terms to enhance their understanding. We also offer detailed definitions and several examples of each choice to help annotators understand how to make decisions. Each entry requires the worker to annotate using a four-point Likert scale. Workers are asked to rate the plausibility of the given question using such scale, where 1 signifies strong agreement and 4 indicates strong disagreement. We consider annotations with a value of 1 or 2 as plausible and those with a value of 3 or 4 as implausible. A snapshot of our annotation instructions, along with a snapshot showing the question released to the worker, are shown in Figure 6 and Figure 7. To ensure comprehension, we require annotators to confirm that they have thoroughly read the instructions by ticking a checkbox before starting the annotation task. We also manually monitor the performance of the annotators throughout the annotation process and provide feedback based on common errors. Spammers or underperforming workers will be disqualified. The overall inter-annotator agreement (IAA) stands at 81% in terms of pairwise agreement, and the Fleiss kappa [20] is 0.56. These statistics are generally comparable to or slightly higher than those of other high-quality dataset construction works [54, 18, 19, 29], which indicates that the annotators are close to achieving a strong internal agreement.
E.3 Expert Verification
@@
Appendix F Task-Oriented Benchmark Comparisons
Table 1 shows a comparison of [Uncaptioned image]Mars with several other datasets, underscoring the unique value of [Uncaptioned image]Mars. In this section, we delve deeper into the differences between various benchmarks for each task, and further elaborates on the distinctive characteristics of [Uncaptioned image]Mars.
F.1 Metaphysical Event Discrimination
In the task of metaphysical event discrimination, we compare [Uncaptioned image]Mars with the discriminative event conceptualization task in AbstractATOMIC and the abstraction detection task in AbsPyramid. Both tasks aim to determine whether a concept feasibly represents an instance within an event (instance abstraction) or the entire event (event abstraction). Despite their similarities to the metaphysical event discrimination task, there are several notable differences. Firstly, none of the previous benchmarks encompass instances of temporal quantifiers, spatial quantifiers, and numerical properties in events, thereby limiting their coverage of instances. This sacrifices a large number of potential changes that could occur within events. Secondly, the concepts in their formulation are disorganized, unlike the increasing abstractive granularity collected in [Uncaptioned image]Mars. Lastly, the primary objective of metaphysical event discrimination is to assess a language model’s ability to discern various abstractions as changes in events, rather than merely evaluating their plausibility in representing instances as concepts.
F.2 Metaphysical Inference Discrimination
In the task of metaphysical inference discrimination, [Uncaptioned image]Mars shares a similar objective with the discriminative triple conceptualization task in AbstractATOMIC. Both tasks evaluate the plausibility of the inference of an abstracted event. However, AbstractATOMIC is limited to featuring social events and, consequently, social inferences. It also only contains nine commonsense relations, and all inference tails are sourced from ATOMIC [54], resulting in very limited semantic coverage. Conversely, [Uncaptioned image]Mars covers a broad range of text events and inferences within various contexts, thanks to the robust generative ability of ChatGPT. Mars also features inferences when the same event is conceptualized in different ways by abstracting different components. This unique feature provides additional value in studying the transition of inferences caused by varying abstractions or variations as changes in the event.
F.3 Metaphysical Transition Discrimination
To the best of our knowledge, no previous benchmarks have covered similar task objective as the metaphysical transition discrimination task. The most comparable tasks are those related to reasoning with changes in logical reasoning or planning, which aim to determine the next necessary step to achieve a goal. This is somewhat akin to inferring the required change in an event to make a metaphysical inference plausible. However, previous works primarily rely on game datasets or feature only a limited number of handcrafted examples, which restricts their effectiveness in evaluating a reasoner’s ability to generalize and understand the consequences of changes across broad domains. Mars addresses this limitation by incorporating a variety of events sourced from Wikitext and Bookcorpus. Previous works also tend to focus solely on selecting the next step from a finite set of possible steps, rather than in an open-ended generative manner. Mars, on the other hand, utilizes ChatGPT to gather additional changes that drive transitions, making it significantly more challenging to reason with transitions in an open-world setting. This approach, however, promotes the development of a generalizable agent with System II reasoning capabilities.

### v2

@@ B.1.4 Metaphysical Transition Generation
Component Type Identified Modified
ME. MI. MT. #Avg. ME. MI. MT. #Avg.
Subject 4,376 3,907 3,507 1.116 3,106 2,950 2,591 1.094
Verb 9,874 8,856 8,061 3.647 4,408 4,146 3,760 3.457
Object 12,645 11,302 9,986 1.760 5,949 5,494 4,865 1.703
Temporal Quantifier 3,003 2,560 2,288 0.472 1,394 1,253 1,110 0.435
Spatial Quantifier 3,866 3,741 3,301 0.459 2,064 1,979 1,718 0.476
Numerical Properties 5,619 4,932 4,355 0.652 3,570 3,353 2,920 0.612
Sub-events 419 385 326 0.040 425 402 332 0.037
Total 39,802 35,683 31,824 8.146 20,916 19,577 17,296 7.814
B.2 Main Evaluations on [Uncaptioned image]Mars
To evaluate LLMs on three tasks in [Uncaptioned image]Mars, we show our evaluating prompts in zero-shot scenario in Table 6. Note that we are aware that LLMs may not be familiar with the word “metaphysical.” Therefore, we also experimented with replacing the word with “implausible,” and the best performances from both types of prompts are reported. These models are consistent across all models’ evaluations for fair comparison.
For few-shot evaluations, few shot examples are added after task descriptions and before the prompted test entry. The exemplars are randomly sampled for each different test entry. For COT prompting, we specifically ask LLMs to “think step by step and generate a short rationale to support your reasoning.” Then, we ask it to give an answer based on its generated rationale. The sampling temperature $\tau$ is set to 0.1 by default, and 5 COT responses are sampled with $\tau$ set to 0.7 in the SC-COT setting.
B.3 Leveraging Open-sourced LLM for Benchmark Curation
In this paper, we use proprietary LLMs and human annotation for data construction, which can be expensive and labor-intensive. However, this approach serves the best pursuit of data quality, which is crucial for an evaluation benchmark. Prior to our data collection, we tested a wide variety of LLMs, and ChatGPT outperformed almost all of them. Therefore, we opted to use it for data construction. Nevertheless, with the recent advancements in state-of-the-art LLMs, we have found that meta-llama/Llama-3.1-405B-Instruct and GPT-4o also achieve satisfactory performance within our data collection framework. We sampled 500 original data entries and employed similar prompts and data collection processes to gather metaphysical reasoning evaluation data entries. We then asked expert annotators to rate the plausibility of the obtained data. The results are shown in Table 4. We observe that LLAMA3.1-405B can achieve comparable performance to ChatGPT in terms of plausible data (evaluation data that reflects reality rather than metaphysics, similar to the majority vote results in Table 2) and expert acceptance rates. Additionally, we find that GPT-4o can even improve the data collection process, resulting in higher quality data. Thus, we believe this represents a compromise between data quality, reproducibility, and cost. It would also be feasible for data collectors to use LLAMA3.1 in the future for collecting metaphysical data, although leveraging proprietary LLMs can be more reliable to some extent.
B.4 Additional Statistics on [Uncaptioned image]Mars
@@
Appendix C Implementation Details
@@
For all experiments, we use the Huggingface [^1] https://huggingface.co/ Library Wolf et al. (2020) to build all models. For each LLM, we conduct experiments with both its instruction fine-tuned version (if any) and the original version. The one achieving higher performances will be included in the reported results. For LLaMa2, the model code is meta-llama/Llama-2-7b/13b/70b(-chat)-hf. For LLaMa3, the model code is meta-llama/Meta-Llama-3-8B/70B(-Instruct). For Mistral, we use mistralai/ Mistral-7B(-Instruct)-v0.3.
@@
All experiments are conducted on eight NVIDIA-V100 (32G) GPUs, with 8E disk space, 48 CPU cores, and 1T memory. Each experiment is repeated three times with different random seeds, and the average performances are reported. The variance across all experiments remains below 0.08, which is considered extremely small. Due to space constraints, we omit reporting this variance.
C.1 Main Evaluations on [Uncaptioned image]Mars
@@
To evaluate PTLMs in a zero-shot manner, we adopt the evaluation pipeline used for zero-shot question answering Ma et al. (2021); Wang et al. (2023a). Specifically, we convert each discrimination data entry into two declarative statements, which serve as natural language assertions corresponding to ‘yes” or “no” options. For instance, when determining whether an event is metaphysical, we generate two assertions: “The event <EVENT> is metaphysical as it’s unlikely to occur in reality,” and “The event <EVENT> is not metaphysical; it’s plausible in reality.” The models are then tasked with computing the loss of each assertion. The assertion with the lowest loss is considered as the model’s prediction. This approach allows any PTLM to be evaluated under classification tasks with an arbitrary number of options or even type classification based on a single assertion. We use the open code library [^3] https://github.com/Mayer123/HyKAS-CSKG as our code base and follow the default hyperparameter settings. For VERA, we follow the exact same implementation [^4] https://github.com/liujch1998/vera Liu et al. (2023a). The accessed backbone model is liujch1998/vera, and all other hyperparameter settings follow the default implementation.
For fine-tuning PTLMs, we connect each PTLM backbone with five fully connected classification layers. The entire model is then fine-tuned using a classification objective with cross-entropy loss. We employ a default setting of a learning rate of 5e-6 and a batch size of 64. The models are optimized using an AdamW optimizer Loshchilov and Hutter (2019), with the model’s performance evaluated every 50 steps. We set the maximum sequence lengths for the tokenizers to 70 for all three discriminative subtasks. Early stopping is also implemented to select the best checkpoint when the highest validation accuracy is achieved. To ensure convergence, we train all models with five epochs.
@@
For fine-tuning LLMs, we use LoRA for fine-tuning, and the LoRA rank and $\alpha$ are set to 16 and 32, respectively. We adopt the open code library from LlamaFactory [^5] https://github.com/hiyouga/LLaMA-Factory Zheng et al. (2024) for model training and evaluation. We similarly use an Adam Kingma and Ba (2015) optimizer with a learning rate of 5e-5 and a batch size of 8. The maximum sequence length for the tokenizer is set at 300. All models are fine-tuned over three epochs, selecting the checkpoint with the highest accuracy on the validation set.
Finally, for evaluating proprietary LLMs, such as ChatGPT and GPT4, we similarly prompt them as with open LLMs. Detailed prompts are explained in Appendix B.2.
We also include full evaluation results (with more baselines and models included) in Table 8. Specifically, for RAG Gao et al. (2023), we reformulate the traditional paradigm of retrieval-augmented generation for our task by asking an LLM to first identify important concepts from the evaluation data entry, retrieve relevant knowledge from an abstract knowledge base containing information about the concepts, and merge them into the evaluation prompt for making the final prediction on metaphysical reasoning tasks. This approach aligns with the design of our Mars benchmark and provides insights into which method offers more benefits when comparing retrieval to fine-tuning conceptual knowledge into LLMs.
For Multi-Agent Calibration, we adopt the multi-agent deliberation design from Yang et al. (2024), which is a multi-agent confidence calibration system for multiple-choice QA. In this setting, we set up two LLMs. The first LLM generates the initial chain-of-thought response and prediction for each task. The second LLM is prompted with the first LLM’s chain-of-thought response and is asked to analyze the differences. Its reasoning rationale regarding these differences, particularly in the metaphysical realm, is then provided as feedback to the first LLM. The first LLM incorporates this feedback and is asked to regenerate the chain-of-thought rationale and final prediction. This loop continues until the second LLM agrees with the first LLM.
For Self-Reflection Pan et al. (2024), we adopt a straightforward approach to rectify LLM errors by using feedback provided by the LLM itself (self-reflection). In this setting, we first ask an LLM to generate a chain-of-thought response explaining the rationale behind a given metaphysical data entry. We then prompt it for a new round, deliberately asking it to analyze the correctness of its rationale and answer. This feedback is merged back into the original prompt and first response to generate a refined response after self-reflection.
C.2 Improving Metaphysical Reasoning via Transferring from Conceptualization Taxonomy
@@
CANDLE’s data is primarily divided into two sections. The first section comprises conceptualizations of instances or events, which can be reformatted into metaphysical event discrimination. Each data entry in CANDLE represents a conceptualization of an abstracted instance within an event or the abstraction of an entire event. Following our definition in Section 3, we interpret each conceptualization as a change in the event. For each data entry, replacing the original instance with its conceptualization forms a plausible change that could occur in reality. Subsequently, we randomly select negative conceptualizations for an event from conceptualizations of other events that do not share any common words with the anchor event. These negative conceptualizations form metaphysical events. Three models are then pre-trained on four million events, with a balanced ratio of plausible events and metaphysical events. The hyperparameters for fine-tuning all models remain consistent with the implementation details described above in Appendix C.1.
@@
Appendix D Annotation Details
D.1 Worker Selection Protocol
@@
D.2 Annotation Interface
For each task, we provide workers with comprehensive task explanations in layman’s terms to enhance their understanding. We also offer detailed definitions and several examples of each choice to help annotators understand how to make decisions. Each entry requires the worker to annotate using a four-point Likert scale. Workers are asked to rate the plausibility of the given question using such scale, where 1 signifies strong agreement and 4 indicates strong disagreement. We consider annotations with a value of 1 or 2 as plausible and those with a value of 3 or 4 as implausible. A snapshot of our annotation instructions, along with a snapshot showing the question released to the worker, are shown in Figure 6 and Figure 7. To ensure comprehension, we require annotators to confirm that they have thoroughly read the instructions by ticking a checkbox before starting the annotation task. We also manually monitor the performance of the annotators throughout the annotation process and provide feedback based on common errors. Spammers or underperforming workers will be disqualified. The overall inter-annotator agreement (IAA) stands at 81% in terms of pairwise agreement, and the Fleiss kappa Fleiss (1971) is 0.56. These statistics are generally comparable to or slightly higher than those of other high-quality dataset construction works Sap et al. (2019); Fang et al. (2021a, b); Hwang et al. (2021); Li et al. (2025); Bai et al. (2023), which indicates that the annotators are close to achieving a strong internal agreement.
D.3 Expert Verification
@@
Appendix E Additional Experiments and Analysis
In this section, we include additional analytical experiments to provide better support for our claims in Mars.
E.1 Multi-task Fine-tuning on [Uncaptioned image]Mars
E.1.1 Setup
To achieve conscious processing, an ideal language model should be capable of performing three tasks uniformly and sequentially. However, fine-tuning each task separately contradicts this objective, as it results in a model that can only perform one task after one training. Therefore, in this section, we investigate the possibility of enabling a language model to master all tasks simultaneously through multitask fine-tuning. Given that all three tasks are binary classification tasks, we adopt a straightforward approach. The language model is trained using a randomly shuffled combination of training data from all three tasks. This anticipates that the model will learn all tasks collectively. The best checkpoint is chosen based on achieving the highest accuracy on the validation sets of all three tasks. After training, the model performance is evaluated separately on the testing sets of each task. All training details remain consistent with those explained in the Appendix C.1.
E.1.2 Results and Analysis
E.2 Few-shot Fine-tuning on [Uncaptioned image]Mars
E.2.1 Setup
From the main evaluation results in Table 2, it is evident that fine-tuning consistently enhances the performance of all models on [Uncaptioned image]Mars. In this section, we delve deeper into the impact of fine-tuning in a few-shot setting, with the aim of analyzing the performance of models trained with limited data. More specifically, we aim to examine how models perform with varying sizes of training data. This will enable us to determine whether collecting more data invariably benefits fine-tuning, thereby leading to the development of more robust metaphysical reasoners. To achieve this, we sample the training data for each task in a progressively increasing ratio of 0.2, 0.4, 0.6, 0.8, and 1.0, and use each sampled training data to fine-tune LLMs for each task individually. The models are then evaluated on the complete validation sets to select the optimal checkpoint, and on the full testing set for performance assessment. All fine-tuning parameters remain consistent across all models, as detailed in Appendix C.1.
Data Split Evaluation Method Event-ACC Inference-ACC Transition-ACC
MARS Zero-shot 53.90 51.20 49.41
MARS Few-shot 49.85 51.47 48.88
MARS-Claude Zero-shot 54.50 54.00 53.50
MARS-Claude Few-shot 56.00 55.50 54.00
MARS-LLAMA3.1 Zero-shot 52.00 56.50 56.50
MARS-LLAMA3.1 Few-shot 55.50 57.50 58.50
Table 7: Evaluation results (%) of GPT-4o on Mars constructed with different backbone LLMs.
E.2.2 Results and Analysis
The results are reported in Table 10. From these results, we observe that training the model with a few-shot training data sample generally has a negative impact across all tasks in [Uncaptioned image]Mars. However, this impact is not significant, and on rare occasions, the sampled training data even leads to superior results compared to training on the full sets. When the training data is reduced to different ratios (80%, 60%, 40%, and 20%), the performance of the models is not significantly affected. This suggests that the models are capable of learning from a small amount of training data and that performance is not significantly influenced by the size of the training data. In other words, annotating more data for training does not necessarily result in better performance, indicating that our task cannot be simply resolved by increasing training data. Future research can explore more advanced reasoning paradigms or training methods to further enhance the capabilities of LLMs in metaphysical reasoning.
E.3 Fine-tuned PTLMs vs. Fine-tuned LLMs
To validate the reason why fine-tuned PTLMs perform better than fine-tuned LLMs, we first hypothesis that PTLMs have a faster convergence rate to the training data due to their smaller number of parameters and fully fine-tuned paradigm (compared to LoRA when fine-tuning LLMs). This results in better fine-tuned performance than LLMs. Although LLMs have lower performance, they exhibit stronger generalizability to other tasks. We fine-tune a DeBERTa-v3 model with 25% and 50% of the training data and observed their performance in Table 10. From the results, we observe that when we reduce the training data for PTLMs, they are hardly comparable to fine-tuned LLMs. However, the last 50% of randomly sampled data brought significant improvements. While we cannot determine the exact reason due to the black box nature of these language models, we believe that PTLMs have a faster rate of fitting into the distribution of the training data or human annotations, resulting in better outcomes on human-annotated evaluation sets. LLMs are more likely to learn how to make correct inferences rather than simply fitting the data. Another possible reason is that we use LoRA to fine-tune LLMs due to limited computational resources; fully fine-tuning LLMs might further enhance their performance.
E.4 Inherent Bias in Mars Construction
One concern regarding the Mars benchmark is the potential bias introduced by using GPT-series models, specifically ChatGPT, for dataset construction. Our approach to constructing Mars was guided by the need to balance scalability with quality. In pilot studies evaluating metaphysical reasoning across various models, GPT-series models consistently demonstrated the highest levels of creativity and reliability. Based on these findings, we selected GPT as the primary backbone for data generation. Constructing Mars, however, required extensive manual annotation, as LLMs often fail to provide accurate labels for complex reasoning tasks. This manual verification process made it impractical to create multiple versions of Mars using different backbone LLMs due to expensive human labors required. Thus, to address concerns about potential biases arising from reliance on ChatGPT, we conducted additional experiments by constructing two smaller versions of the Mars benchmark. These alternative benchmarks utilized data generated from two different LLMs, Claude-3.5-sonnet Anthropic (2024) and LLAMA 3.1-70B Dubey et al. (2024), in each step, to obtain 200 evaluation data entries per task in Mars. All samples underwent expert annotation to collect ground-truth labels. We then evaluate GPT-4’s zero-shot and few-shot performance on these alternative benchmarks alongside the original Mars.
The results are shown in Table 7. We observe that using different LLMs as backbones for Mars construction results in similar performance by GPT-4 across zero-shot and few-shot settings. Overall, the difficulty of the Mars benchmark remains robust and consistent, irrespective of the backbone LLM used during dataset generation. These experiments demonstrate that the reliance on ChatGPT for the original Mars construction does not compromise the benchmark’s validity or difficulty. The results reinforce the reliability of MARS as a comprehensive test of metaphysical reasoning, with its complexity surpassing any potential biases introduced by the specific LLM used in data collection.
E.5 Binary Task Design in Mars
In Mars, all tasks are designed as a binary prediction task to facilitate automated and easy label collection and evaluation. Here, we discuss the reason and some pilot analysis behind such task design by considering other task formulations, including multiple-choice, open-ended generation, and binary evaluation.
Multiple-choice tasks, while structured and amenable to automated evaluation, posed significant challenges in collecting high-quality negative (distractor) options. Relying on human annotators to create distractors proved labor-intensive and impractical for scaling, as it required drafting multiple plausible but incorrect options for each question. As a result, we adopted open-ended generation and binary evaluation, ultimately choosing a generate-then-annotate paradigm. This approach involved two stages: first, evaluating the performance of LLMs in generating metaphysical cases during the generation phase; second, annotating the generated cases with binary labels (correct/incorrect).
To complement the binary evaluation results, we also included human annotation results for ChatGPT’s performance in generating metaphysical data, as indicated in the Majority row of Table 2, which can be regarded as following a generative task paradigm. The results demonstrate that, even when the task is framed as a generation task, ChatGPT struggles with metaphysical reasoning. The low proportion of human-annotated correct generations highlights the difficulty of reasoning about metaphysical changes, regardless of task formulation. While binary evaluation offers clear performance metrics and scalability advantages, the generation task provides complementary insights into the model’s creative and reasoning capabilities. Together, these observations underscore the importance of improving LLMs’ ability to reason about distributional and situational changes, which is crucial for advancing their metaphysical reasoning capabilities.
Appendix F Case Studies
In this section, we present some examples for each of the three tasks in [Uncaptioned image]Mars to help readers better understand our benchmark. The examples are displayed in Table 11. We observe that examples in [Uncaptioned image]Mars typically require careful reasoning and consideration of the plausibility of occurrences in reality or the metaphysical realm to make the correct discrimination.
Methods Backbone Event Inference Transition
Acc AUC Ma-F1 Acc AUC Ma-F1 Acc AUC Ma-F1
Random - 50.00 - 49.56 50.00 - 49.56 50.00 - 49.56
Majority - 60.98 - 37.99 58.56 - 36.93 50.25 - 33.37
GPT2-XL 1.5B 38.62 51.12 27.93 44.40 51.88 31.45 49.92 48.35 48.09
CAR 435M 54.63 49.34 49.96 48.33 42.85 41.93 52.97 35.05 46.94
CANDLE 435M 51.90 49.12 50.30 46.77 44.03 38.48 53.49 34.95 47.95
VERA 11B 51.82 50.48 48.52 60.97 62.54 59.09 61.31 66.32 61.17
GPT2-XL 1.5B 46.68 47.63 46.96 43.70 44.22 30.41 44.57 45.03 45.89
VERA 11B 61.95 61.43 60.81 63.90 66.93 70.84 71.75 74.57 73.27
Meta-LLaMa-2-13B 51.50 - 49.48 50.81 - 50.57 50.81 - 50.80
Meta-LLaMa-2-70B 52.40 - 49.03 56.13 - 46.81 48.45 - 48.34
Meta-LLaMa-3-8B 50.62 - 49.12 51.33 - 50.98 51.95 - 51.07
Meta-LLaMa-3-70B 57.41 - 50.59 63.40 - 61.82 60.15 - 60.01
Meta-LLaMa-3.1-8B 51.01 - 50.27 52.13 - 51.29 52.35 - 52.09
Meta-LLaMa-3.1-70B 59.22 - 52.08 63.61 - 61.90 61.28 - 61.03
+RAG 61.21 - 54.51 66.38 - 65.90 61.53 - 61.22
+Multi-Agent 56.12 - 51.08 65.06 - 65.01 62.54 - 62.19
+Self-reflection 57.94 - 53.17 63.91 - 63.51 60.92 - 60.77
Meta-LLaMa-3.1-405B 59.22 - 52.08 63.61 - 61.90 61.28 - 61.03
Gemma-2-9B 56.88 - 48.53 51.83 - 51.76 49.41 - 45.01
Falcon-7B 54.32 - 49.51 51.77 - 50.30 50.42 - 49.02
Falcon-40B 52.35 - 50.36 49.67 - 49.38 50.27 - 50.22
Mistral-7B 49.90 - 48.94 50.23 - 50.06 51.75 - 51.75
Gemma-2-9B 61.23 61.25 60.28 69.24 70.76 69.00 73.30 76.91 69.18
Mistral-7B 60.35 60.77 60.07 66.91 70.06 65.95 71.87 75.47 68.53
LLM (API) ChatGPT 51.00 - 50.35 61.35 - 57.63 60.40 - 60.12
ChatGPT (5-shots) 53.61 - 53.28 58.05 - 57.42 62.40 - 59.35
ChatGPT (COT) 53.20 - 52.61 50.40 - 50.32 49.95 - 49.83
ChatGPT (SC-COT) 53.98 - 53.47 52.47 - 51.99 51.25 - 51.13
GPT4 53.90 - 53.45 51.20 - 50.95 49.41 - 49.33
GPT4 (5-shots) 49.85 - 49.58 51.47 - 51.30 48.88 - 48.73
GPT4 (COT) 51.28 - 50.73 51.49 - 51.35 47.62 - 47.58
GPT4 (SC-COT) 51.97 - 51.26 52.05 - 52.27 48.24 - 48.11
GPT-4o-mini 57.94 - 57.91 53.84 - 53.53 48.06 - 48.06
+RAG 59.99 - 59.97 54.54 - 54.21 49.39 - 49.19
+Multi-Agent 54.21 - 53.17 52.76 - 52.26 46.94 - 46.70
+Self-reflection 56.89 - 55.21 53.22 - 53.20 48.51 - 48.45
Table 8: Full evaluation results (%) of various language models on the testing sets of Mars. The best performances within each method are underlined and the best among all methods are bold-faced.